### v1

# Information-Theoretic Reward Decomposition for Generalizable RLHF

Abstract
A generalizable reward model is crucial in Reinforcement Learning from Human Feedback (RLHF) as it enables correctly evaluating unseen prompt-response pairs. However, existing reward models lack this ability, as they are typically trained by increasing the reward gap between chosen and rejected responses, while overlooking the prompts that the responses are conditioned on. Consequently, when the trained reward model is evaluated on prompt-response pairs that lie outside the data distribution, neglecting the effect of prompts may result in poor generalization of the reward model. To address this issue, we decompose the reward value into two independent components: prompt-free reward and prompt-related reward. Prompt-free reward represents the evaluation that is determined only by responses, while the prompt-related reward reflects the reward that derives from both the prompt and the response. We extract these two components from an information-theoretic perspective, which requires no extra models. Subsequently, we propose a new reward learning algorithm by prioritizing data samples based on their prompt-free reward values. Through toy examples, we demonstrate that the extracted prompt-free and prompt-related rewards effectively characterize two parts of the reward model. Further, standard evaluations show that our method improves both the alignment performance and the generalization capability of the reward model.
Machine Learning, ICML
Figure 1: Left: reward gaps calculated with corresponding prompt and randomly sampled prompts using QRM-Llama3-8B 11footnotemark: 100footnotetext: xx on two different datasets that were used for training. When calculating with other prompts, the curves show the mean and the std. Right: illustrative failure case where the reward gap overly depends on the responses. Solid lines represent corresponding prompt-response pairs (used in training), while dashed lines represent non-corresponding pairs (unseen during training). Since the reward gap overly depends on the responses, it generalizes poorly to novel prompt-response pairs constructed even with seen but non-corresponding prompts.
1 Introduction
Reinforcement Learning from Human Feedback (RLHF) is an effective approach for Large Language Models (LLMs) alignment (Christiano et al., 2017; Bai et al., 2022). Within a wide range of RLHF methods, reward learning plays a pivotal role. These methods typically first train a reward model on a static dataset and then leverage it to do Reinforcement Learning (RL) (Ouyang et al., 2022; Dong et al., 2024). Compared with methods that are free of using reward models (Rafailov et al., 2024b; Zhao et al., 2023), the advantage of such methods is their capacity to leverage the generalization capability of the reward model to evaluate outside-of-distribution prompt-response pairs. These prompt-response pairs with generated rewards can be used to further improve the LLM’s performance (Stiennon et al., 2020; Ziegler et al., 2019).
Clearly, learning a generalizable reward model is central to this scenario. However, we found that standard reward training does not guarantee sufficient generalization capability. In reward model training, the primary goal is typically better distinguishing between chosen and rejected responses. To achieve this, the reward model does not necessarily require consideration of the corresponding prompt. Taking reward learning based on Bradley-Terry (BT) model as an example. Since the potential response space is vastly larger than the dataset size, different data samples typically contain distinct response pairs. As long as the reward gap within each pair increases, the training loss will decrease effectively. This can occur even if the reward model only takes the responses as the input. For the reward model that separately calculates rewards for chosen and rejected prompt-response pairs, the situation becomes similar as the inputs to the reward model only differ starting from the response.
Perhaps surprisingly, such a phenomenon indeed appears in current reward models, even those that achieve SOTA performance on benchmarks. As shown in Fig. 1 (left), after replacing the corresponding prompt with other prompts, the reward gaps still center around their original values. This issue, where responses dominate the reward gap, does not affect training but leads to catastrophic results when evaluating novel prompt-response pairs. The illustrative example in Fig. 1 (right) shows this. When considering each prompt-response pair separately within the training dataset, the reward model appears to perform perfectly. However, when querying preferences after replacing the original prompt with other even in-distribution prompts (which are also meaningful queries), the reward model can yield significantly inaccurate or even reversed preferences. This generalization issue will become more pronounced when dealing with unseen prompt-response pairs encountered during evaluation. All of these highlight the need to distinguish two components of the reward value: one part is the value determined solely by the response, and the other is the value that can only be derived by simultaneously considering both the prompt and the response. We refer to the former as prompt-free reward and the latter as prompt-related reward.
To address this, we propose a novel method of decomposition to extract these two components from an information-theoretic perspective, without requiring extra models. After that, we use the extracted prompt-free reward to guide the reward learning process, prioritizing training samples based on their prompt-free reward gap values. We verify our method through several toy examples and standard evaluations based on commonly used datasets and base models. In toy examples, the extracted prompt-free and prompt-related reward gaps characterize the reward model’s preference bias about response-only features and the generalizable preference information, respectively. Moreover, in standard experiments with common datasets, the reward model trained with our method outperforms strong baselines. These experiments show that considering both prompt-free and prompt-related rewards during training enhances the alignment performance and generalization capabilities of the reward model. footnotetext: 1https://huggingface.co/nicolinho/QRM-Llama3-8B.
2 Preliminaries
Standard preference learning assumes the existence of the preference oracle which determines $\mathbb{P}(y_{w}\succ y_{l}|x)$, where the probability of response $y_{w}$ is more preferred than $y_{l}$ conditioned on the prompt $x$. Given a preference dataset $D=\{(x,y_{w},y_{l})_{i}\}_{i=1}^{N}$ where the prompt-response tuples $(x,y_{w},y_{l})$ are generated following $\mathbb{P}(y_{w}\succ y_{l}|x)$, our objective is to estimate the preference oracle from it.
There exist several methods to model the preference oracle (Luce, 1959), among which the Bradley-Terry (BT) model (Bradley & Terry, 1952) is the most widely used. The BT model further assumes that the preference oracle can be represented as a reward model $r:(x,y)\to\mathbb{R}$ that satisfies:
$\mathbb{P}_{r}(y_{w}\succ y_{l}|x)=\frac{\exp(r(x,y_{w}))}{\exp(r(x,y_{w}))+\exp(r(x,y_{l}))}.$ (1)
Under this assumption, standard methods leverage a parameterized reward model $r_{\theta}$ to perform maximum likelihood estimation on the preference dataset (Ouyang et al., 2022; Stiennon et al., 2020) via the following objective:
$\max\mathbb{E}_{(x,y_{w},y_{l})\sim D}[\log\mathbb{P}_{r_{\theta}}(y_{w}\succ y_{l}|x)].$ (2)
In this work, we only consider the preference learning based on the BT model, owing to its widespread use and strong performance in the field of RLHF. The parameterized reward model $r_{\theta}$ can be implemented in various ways (Touvron et al., 2023; Dong et al., 2024; Wang et al., 2024; Ivison et al., 2023). Among these, the most commonly used one is the ‘Sequence Classifier’ (Lambert et al., 2024; Ouyang et al., 2022), which builds the reward model on top of an LLM backbone. By projecting the feature representation of the prompt-response sequence to a scalar value using a simple network (e.g., single linear layer), such a reward model effectively leverages the prior knowledge embedded in the LLM backbone to evaluate prompt-response pairs. In this work, we focus on reward models within this type.
3 Methods
In this section, we first present a specific form of decomposition that divides the reward value into a free reward and a related reward. We then formalize the problem of such decomposition from a given reward model and propose a solution from an MI perspective. We also demonstrate that the solution can be effectively obtained without extra models.
3.1 Prompt-free Reward & Prompt-related Reward
We decompose the reward value $r_{\theta}(x,y)$ into two separate parts: the prompt-free reward and the prompt-related reward. The prompt-free reward is only determined by the response and can be regarded as the overall evaluation of the response. Specifically, the prompt-free reward ($r_{2}$) satisfies Eq. (3). Once the response is given, the prompt-free reward remains unaffected by any specific prompt, as
$\forall x_{1},x_{2},y~{}~{}\in D,~{}r_{2}(x_{1},y)=r_{2}(x_{2},y).$ (3)
In contrast, as the other part of $r_{\theta}(x,y)$ besides the prompt-free reward, prompt-related reward ($r_{1}$) varies among different prompts. Such a reward can be determined only when both the prompt and the response are given.
Given a reward model $r_{\theta}$ during training and a prompt-response tuple $(x,y_{1},y_{2})$ sampled from the dataset, our goal is to identify the prompt-free reward in $r_{\theta}(x,y_{1})$ and $r_{\theta}(x,y_{2})$ and leverage it to guide reward learning. We consider decomposing $r_{\theta}(x,y)$ into the additive form:
$r_{\theta}(x,y)=r_{1}(x,y)+r_{2}(x,y).$ (4)
Examining the dataset holistically, $r_{2}(x,y)$ demonstrates clear randomness, with the data distribution influencing its value. Specifically, for a given $(x,y)$ pair, if the prompts related to $y$ are diverse, there is little connection between $r_{2}(x,y)$ and $r_{\theta}(x,y)$, since $r_{2}$ remains unchanged across different prompts, whereas $r_{\theta}$ changes accordingly. Conversely, if the related prompts exhibit little variety, $r_{2}(x,y)$ can be closely related to $r_{\theta}(x,y)$, as the variation in $r_{\theta}(x,y)$ is also limited. This motivates us to define $r_{2}$ by considering the randomness in the data distribution, treating $r_{2}$ as the joint product of $r_{\theta}$ and the data distribution.
To obtain the overall evaluation of $r_{\theta}$ on $y$, one may consider simple transformations like directly taking the expectation of $r_{\theta}(x,y)$ over the prompt distribution, i.e. $r_{2}(x,y)=\mathbb{E}_{x^{\prime}\sim P(X|Y=y)}[r_{\theta}(x^{\prime},y)]$. Although $r_{2}(x,y)$ defined in this way is independent of $x$, such a definition can be inconsistent with ‘prompt-free reward’ from the preference perspective. To illustrate this, consider two responses, $y_{1}$ and $y_{2}$. If $r_{\theta}(x,y_{1})$ is significantly larger than $r_{\theta}(x,y_{2})$ under a single prompt, then even if $r_{\theta}(x,y_{2})$ is slightly larger than $r_{\theta}(x,y_{1})$ under all other prompts, $r_{2}(x,y_{1})$ will be much larger than $r_{2}(x,y_{2})$. Such a prompt-free reward strongly favors $y_{1}$, despite $r_{\theta}$ overall preferring $y_{2}$. A detailed numerical example is given in Appendix B. This indicates that a carefully designed objective is required to induce the prompt-free reward.
3.2 Solving Prompt-free Reward via a MI Objective
Although directly solving $r_{2}(x,y)$ by marginalizing $r_{\theta}(x,y)$ is challenging, we can extract $r_{2}(x,y)$ from $r_{\theta}(x,y)$ via a preference perspective. For simplicity, we have notations
$\Delta r(x,y_{1},y_{2})$ $\coloneqq r(x,y_{1})-r(x,y_{2})$ (5)
$\sigma(\Delta r(x,y_{1},y_{2}))$ $\coloneqq 1/\big{(}1+\exp(-\Delta r(x,y_{1},y_{2}))\big{)}.$ (6)
Since we focus on preference learning based on the BT model, it is sufficient to study the reward gap ($\Delta r$) rather the reward itself, as the reward gap decides the preference label. Consequently, we decompose the reward gap as
$\Delta r_{\theta}(x,y_{1},y_{2})=\Delta r_{1}(x,y_{1},y_{2})+\Delta r_{2}(x,y_{1},y_{2}).$ (7)
Meanwhile, since both $r_{1}$ and $r_{2}$ must be applicable to all $(x,y_{1},y_{2})$, it is essential to take into account the entire dataset when considered from the preference perspective. The preferences of $r_{1}$, $r_{2}$, and $r_{\theta}$ over the entire dataset are inherently random and can be characterized by the following random variables:
$\tilde{Z}$ $=\text{Ber}\big{(}\mathbb{P}_{r_{2}}(Y_{1}\succ Y_{2}|X)\big{)}=\text{Ber}\big{(}\sigma(\Delta r_{2}(X,Y_{1},Y_{2})\big{)},$ (8)
$Z$ $=\text{Ber}\big{(}\mathbb{P}_{r_{1}}(Y_{1}\succ Y_{2}|X)\big{)}=\text{Ber}\big{(}\sigma(\Delta r_{1}(X,Y_{1},Y_{2})\big{)},$
$\tilde{W}$ $=\text{Ber}(\mathbb{E}_{x\sim P(X|Y_{1},Y_{2})}[\mathbb{P}_{r_{\theta}}(Y_{1}\succ Y_{2}|x)])$
$=\text{Ber}(\mathbb{E}_{x\sim P(X|Y_{1},Y_{2})}[\sigma(\Delta r_{\theta}(x,Y_{1},Y_{2}))]),$
$W$ $=\text{Ber}\big{(}\mathbb{P}_{r_{\theta}}(Y_{1}\succ Y_{2}|X)\big{)}=\text{Ber}\big{(}\sigma(\Delta r_{\theta}(X,Y_{1},Y_{2}))\big{)},$
where ‘Ber’ stands for Bernoulli random variable. Among them, $W$, $Z$, and $\tilde{Z}$ are the random preferences of their corresponding rewards. $\tilde{W}$ is the prompt-free random preference of $r_{\theta}$. For all these random variables, the randomness comes from the prompt, responses, and the Bernoulli distribution. We also provide a detailed interpretation of these random variables in Appendix D.1.
Based on the defined Bernoulli random variables, we proceed to characterize the desired $r_{1}$ and $r_{2}$. Recall that $r_{1}$ and $r_{2}$ represent prompt-related and prompt-free rewards, respectively. This means $Z$ should encapsulate only prompt-related preference while $\tilde{Z}$ should only reflect prompt-free preference. Formally, this could be written in the following conditions:
$\begin{cases}\text{MI}(Z~{}\|~{}\tilde{W})=0,\hfill\\ \text{MI}(\tilde{Z}~{}\|~{}\tilde{W})=\text{MI}(\tilde{Z}~{}\|~{}W).\hfill\end{cases}$ (9)
The Venn diagram in Fig. 2 further illustrates these conditions. The first condition eliminates all information in $Z$ that’s related to prompt-free preference, while the second condition ensures that the information contained in $\tilde{Z}$ does not exceed the prompt-free preference.
Figure 2: Sub-figure (a) illustrates the information in $W$ and $\tilde{W}$. Sub-figures (b) and (c) characterize the desired and undesired pairs of $Z$ and $\tilde{Z}$, respectively.
Although these conditions effectively constrain the information in $Z$ and $\tilde{Z}$, a trivial solution exists. Assume $\Delta r_{\theta}(x,y)$ is bounded for any prompt $x$ and response $y$. Consider an ill-formed $\Delta r_{1}$ such that for any prompt $x$ and response $y$, $\Delta r_{1}(x,y)=+\infty$. Then by the definition that $\Delta r_{\theta}(x,y)=\Delta r_{1}(x,y)+\Delta r_{2}(x,y)$, $\Delta r_{2}(x,y)=-\infty$ for any prompt $x$ and response $y$. In this case, $Z$ will always be $1$, and $\tilde{Z}$ will always be 00. This makes $\text{MI}(Z\|\tilde{W})=\text{MI}(\tilde{Z}\|\tilde{W})=\text{MI}(\tilde{Z}\|W)=0$ since $Z$ and $\tilde{Z}$ are both constants. Although these ill-formed $r_{1}$ and $r_{2}$ satisfy the conditions in Eq. (9), they provide no information of prompt-related or prompt-free preferences. This ill-formed example indicates that the MI condition in Eq. (9) is insufficient to constrain the information in $Z$ or $\tilde{Z}$ themselves. When $Z$ and $\tilde{Z}$ become constants, the information within them is lost, along with their MI with other random variables. To avoid this, we optimize the following constrained objective to derive $r_{1}$ and $r_{2}$ instead:
$\max_{r_{1}}\quad H(Z),\quad\text{s.t.}\quad\left\{\begin{aligned} \text{MI}(Z\|\tilde{W})&=0\\ \text{MI}(\tilde{Z}\|\tilde{W})&=\text{MI}(\tilde{Z}\|W).\end{aligned}\right.$ (10)
In addition to the previous conditions, we maximize the information within $Z$ to ensure a diverse prompt-related preference, considering that $Z$ represents the random prompt-related preference across the entire dataset distribution. This constrained optimization problem is difficult to solve directly, and direct solutions may require additional parameterized reward models. However, we provide an efficient algorithm that obtains $\Delta r_{1}(x,y_{1},y_{2})$ and $\Delta r_{2}(x,y_{1},y_{2})$ without additional reward model. To see this, we first note that the second constraint is satisfied with a specific structure of $r_{2}$. Formally, the following theorem holds:
Theorem 1.
When the value of $r_{2}$ depends only on the response, i.e. $r_{2}(x,y)=r_{2}(y)$, MI$(\tilde{Z}\|\tilde{W})$ $=$ MI$(\tilde{Z}\|W)$.
The proof is given in Appendix D.2. Intuitively, Theorem 1 holds since such $\tilde{Z}$ inherently does not contain any information from the prompt. Consequently, removing the prompt-related information from $W$ to $\tilde{W}$ does not reduce the MI term. This important property enables us to concentrate on the first constraint and maximize the information in $Z$, as long as $r_{2}$ maintains this simple structure. We continue to characterize the optimal solution of Eq. (10) in the following theorem.
Theorem 2.
For any bounded $r_{\theta}$ and dataset $(X,Y_{1},Y_{2})$, there exist feasible $r_{1}^{*}$, $r_{2}^{*}$ such that $\forall(y_{1},y_{2})\sim P(Y_{1},Y_{2}),\mathbb{E}_{x\sim P(X|Y_{1}=y_{1},Y_{2}=y_{2})}[\sigma(\Delta r_{1}^{*}(x,y_{1},y_{2}))]=\frac{1}{2}$. Such $r_{1}^{*}$, $r_{2}^{*}$ is the optimal solution to problem (10).
The proof is given in Appendix D.2. Intuitively, this requires that the prompt-free part of $r_{1}^{*}$ could not distinguish between the responses within any given response pair. This aligns with our desiderata, as $r_{1}^{*}$ should only reflect the evaluations considering both the prompt and response. Although neither of the theorems explicitly shows the connection with the preference of $r_{\theta}$, we note that the feasibility condition requires $\Delta r_{1}^{*}(x,y_{1},y_{2})+\Delta r_{2}^{*}(x,y_{1},y_{2})$ to equal $\Delta r_{\theta}^{*}(x,y_{1},y_{2})$, which in turn avoids trivial solution. Theorem 2 also provides an efficient way to obtain $\Delta r_{1}^{*}(x,y_{1},y_{2})$ and $\Delta r_{2}^{*}(x,y_{1},y_{2})$. After replacing $\Delta r_{1}^{*}(x,y_{1},y_{2})$ with $\Delta r_{\theta}(x,y_{1},y_{2})-\Delta r_{2}^{*}(x,y_{1},y_{2})$ in the optimal solution in Theorem 2, $\Delta r_{2}^{*}(x,y_{1},y_{2})$ satisfies that for any $(y_{1},y_{2})\sim P(Y_{1},Y_{2})$:
$\underset{x\sim P(X|Y_{1}=y_{1},Y_{2}=y_{2})}{\mathbb{E}}[\sigma(\Delta r_{\theta}(x,y_{1},y_{2})-\Delta r_{2}^{*}(y_{1},y_{2}))]=\frac{1}{2}.$ (11)
Note that we replace $\Delta r_{2}^{*}(x,y_{1},y_{2})$ with $\Delta r_{2}^{*}(y_{1},y_{2})$ due to Theorem 1. Such $\Delta r_{2}^{*}(y_{1},y_{2})$ represents a consistent shift across different prompts. By subtracting this shift from $\Delta r_{\theta}(x,y_{1},y_{2})$, we can remove the prompt-free preference in it. It is straightforward to prove that the expectation in Eq. (11) decreases monotonically with the increase of $\Delta r_{2}^{*}(y_{1},y_{2})$, as shown in Appendix D.2. Combined with the bounded assumption of $r_{\theta}$, $\Delta r_{2}^{*}(y_{1},y_{2})$ can be obtained via binary search in a finite interval. The searched $\Delta r_{2}^{*}(y_{1},y_{2})$ can be interpreted as the ‘weighted average’ of $\Delta r_{\theta}(x,y_{1},y_{2})$, based on $P(X|Y_{1}=y_{1},Y_{2}=y_{2})$. Due to limited space, we provide the pseudo-code of the binary search process in Appendix C.1 (Alg. 1).
It’s clear that during the binary search, the estimation of the expectation in Eq. (11) requires sampling from $P(X|Y_{1}=y_{1},Y_{2}=y_{2})$. Although sampling from it is generally challenging, we propose utilizing importance sampling by considering the decomposition with Bayes’ rule:
$P(x|y_{1},y_{2})=P(x)P(y_{1},y_{2}|x)~{}/~{}P(y_{1},y_{2}).$ (12)
Here $P(x|y_{1},y_{2})$ is the abbreviation of $P(X=x|Y_{1}=y_{1},Y_{2}=y_{2})$, so as other probabilities. Given a $(y_{1},y_{2})$, we first sample from $P(X)$, the marginal distribution of prompts, and then re-weight the samples using $P(y_{1},y_{2}|x)$. Since this value represents the generation probability of the responses for a specific prompt, it can be efficiently computed in some cases or approximated in others. We list the sampling schemes for different cases in Appendix C.2.
3.3 Guide Reward Learning with Prompt-free Reward
Figure 3: (Illustrative) We characterize training data samples in a 2-dimensional quadrant diagram, with the decomposed reward gaps $\Delta r_{1}$ and $\Delta r_{2}$. (a) shows initial data samples before training. (b) After training, the ideal distribution should be centered on the positive half of the $\Delta r_{1}$-axis, indicating that the preference depends solely on prompt-related information. (c-d) However, in the training process of $\Delta r_{\theta}$, the update following the BT model can only ensure the data points move in at least one of the positive directions of the $\Delta r_{1}$-axis or $\Delta r_{2}$-axis. If the update is based on samples with small $\Delta r_{2}$ (e.g. the left half), their movement upwards or to the right, along with other unupdated samples (e.g. the right half), will cause the distribution to be more centered on the positive $\Delta r_{1}$-axis. On the other hand, if the update is based on all samples, as shown in (d), the movement of samples with large $\Delta r_{2}$ values to the right exacerbates the existing prejudices, causing the distribution to become more centered on the positive $\Delta r_{2}$-axis.
During the training process of $r_{\theta}$, one may consider leveraging the reward gap value $\Delta r_{\theta}(x,y_{1},y_{2})$ to selectively train more on samples with small reward gap (including negative ones). The reason is that, if the reward gap for $(x,y_{1},y_{2})$ satisfies $\Delta r_{\theta}(x,y_{1},y_{2})\approx 0$ or $<0$, it indicates that the current reward model $r_{\theta}$ struggles to distinguish between the preferred and dispreferred responses. Consequently, allocating more training budget to these samples can better align the reward model with human preference. Similar methods have been studied in training LLMs (Lin et al., 2024; Mindermann et al., 2022).
Characterize ideal $r_{\theta}$ with $\Delta r_{1}$ and $\Delta r_{2}$ However, things become different if we consider prompt-related and prompt-free reward separately. Similar to the method discussed above, $\Delta r_{1}$ and $\Delta r_{2}$ are important properties of $r_{\theta}$ and can be leveraged to guide the training of $r_{\theta}$ itself. Specifically, during the learning process of $r_{\theta}$, we expect (i) $r_{\theta}$ to learn more prompt-related preference from the dataset (large prompt-related reward gaps); and (ii) $r_{\theta}$ to learn less prompt-free preference (small absolute values of prompt-free reward gaps). This is because both very large or very small prompt-free reward gaps (i.e., $\Delta r_{2}(x,y_{1},y_{2})\gg 0$ or $\ll 0$) indicate prejudice over the sampled responses, as they do not take the specific prompt into account. This prejudice will not only influence preferences under other prompts but may also introduce spurious reward gaps, e.g., reward gaps induced by properties only related to the responses, such as length bias (Dubois et al., 2024; Singhal et al., 2024). A more detailed discussion of this is given in §4.1.1.
Prioritizing data with small $\Delta r_{2}$ To fulfill the expectations above, we present a straightforward method that prioritizes preference data in training reward models, with the aim of increasing prompt-related reward gaps while constraining the value of prompt-free reward gaps. Specifically, given data samples $\{(x,y_{1},y_{2})_{i}\}_{i=1}^{k}$ in each iteration, we first calculate their prompt-free reward gaps $\{\Delta r_{2}(x,y_{1},y_{2})_{i}\}_{i=1}^{k}$ based on Alg. 1, and then perform update for $r_{\theta}$ on those samples with small prompt-free reward gaps. For samples with large prompt-free gaps, we ignore them in this iteration and reinsert them into the buffer for subsequent updates.
Analysis of prioritization mechanism.  Next, we provide an in-depth analysis for the proposed data prioritization mechanism. The standard update following the BT model only ensures the increase of $\Delta r_{1}(x,y_{w},y_{l})+\Delta r_{2}(x,y_{w},y_{l})$, so we cannot guarantee that either of them will necessarily increase. However, (i) if samples have small prompt-free reward gaps (e.g. $\Delta r_{2}<0$), regardless of whether their prompt-related reward gaps ($\Delta r_{1}$) are large or small, they should be used for updates. This is because an increase in $\Delta r_{1}$ suggests a better mastery of prompt-related preferences, while an increase in their $\Delta r_{2}$ indicates the elimination of existing prejudices, both of which are beneficial. (ii) Conversely, for samples with large $\Delta r_{2}$ (e.g., $\Delta r_{2}\gg 0$), if their $\Delta r_{1}$ are also large, updating is unnecessary since the prejudice led to gradient saturation in the BT model. If their $\Delta r_{1}$ are small, the preferences on them stem mainly from factors unrelated to the prompt (e.g., response length), which are easier for $r_{\theta}$ to learn from. Updating $r_{\theta}$ on them would exacerbate this issue, hindering generalization. Fig. 4 further illustrates this mechanism. Data samples with smaller $\Delta r_{2}$ help achieve a better trade-off between increasing prompt-related reward gaps and constraining prompt-free reward gaps.
Practical Consideration. In practice, we design a binary-cluster mechanism to determine whether the value of $\Delta r_{2}$ is ‘small’ or ‘large’ in data prioritization. Specifically, at each step, we sample a batch of $(x,y_{1},y_{2})$ pairs from the preference dataset and perform one-dimensional binary clustering based on their $\Delta r_{2}$ values. Due to the dynamic nature of $r_{\theta}$ and the noise in data sampling, we do not directly use the obtained decision boundary for determination. Instead, we maintain a dynamic threshold, calculated as the exponential moving average of the boundary, and use this threshold for the determination. Due to limited space, we provide a detailed process in Appendix C.1 (Alg. 2).
It is worth noting that samples with $\Delta r_{2}$ values exceeding the threshold will not be discarded from the dataset. Instead, the reinsertion operation assigns a probability for these samples to be updated the next time they are sampled. Since $r_{\theta}$ dynamically evolves during training, samples that are not prioritized at the moment may contribute significantly to the future $r_{\theta}$. To avoid endlessly cycling through samples with large $\Delta r_{2}$ values, the opportunities for reinsertion are limited. We provide more details of the algorithm in Appendix C.
4 Experiments
In this section, we demonstrate the significance of identifying and utilizing prompt-free rewards to guide reward learning from two perspectives. We first illustrate that, for some manually crafted datasets, the extracted prompt-free reward reflects the reward model’s preference bias and the prioritization effectively aids in reward learning. Subsequently, based on some commonly used open-source preference datasets, we evaluate the trained reward model using direct metrics (reward model accuracy) and indirect metrics (performance of the induced policy). The evaluation results demonstrate that, with the guidance of prompt-free rewards, our method enhances the generalization ability of the learned reward model and further improves the performance of the induced LLM policy after alignment.
4.1 Experiments on Manually Crafted Datasets
In this part, we manually construct datasets with specific characteristics from SHP (Ethayarajh et al., 2022), a commonly used open-source preference dataset. Based on these datasets, we compare the reward models trained with ordinary data and prioritized data. To thoroughly demonstrate their difference during training, we select four equally spaced training steps, sampling a number of data pairs and visualizing them in the same way in Fig. 4. Additionally, we evaluate the final reward model using Reward-Bench (Lambert et al., 2024) to assess its generalization capabilities. We choose LLaMA-3.2-1B-Instruct as the base model due to its lightweight nature and strong performance. For more details of the experiments, we refer to Appendix E.2.
4.1.1 Length-biased Dataset
We construct a length-biased dataset $\mathcal{D}_{\rm bias}$ that contains preference pairs with 80% chosen-longer responses (i.e. $|y_{w}|>|y_{l}|$) and 20% chosen-shorter responses. The details of the construction are given in Appendix E.2. When performing reward learning in $\mathcal{D}_{\rm bias}$, the BT loss for a uniformly sampled data batch can be easily optimized by considering only the lengths of the responses, as the majority of the data exhibits a preference for longer chosen responses. As a result, the reward model can easily overfit to such spurious, prompt-free preferences.
The visualizations during training and the final results on Reward-Bench are shown in Fig. 4 and Table 2. If $r_{\theta}$ is trained with data uniformly sampled from $\mathcal{D}_{\rm bias}$ (Fig. 4, upper part), although its prompt-related reward gap ($\Delta r_{1}$) increases gradually, $r_{\theta}$ rapidly overfit to the length preference. This results in a significantly faster increase in $\Delta r_{2}$ for the chosen-longer pairs, and inevitably leads to a decrease in $\Delta r_{2}$ for the chosen-shorter pairs since they represent opposite prompt-free preferences. As expected, such $\Delta r_{2}$ reveals a clear length preference within $r_{\theta}$, which undermines the model’s ability to discriminate between the responses within chosen-shorter pairs. On the other hand, if we prioritize the data points with smaller $\Delta r_{2}$ values, the data batch will contain more chosen-shorter pairs with smaller $\Delta r_{2}$ values before $r_{\theta}$ heavily overfits to the length preference. According to Fig. 4 (lower part), the data samples are well-centered around the positive $\Delta r_{1}$-axis during training. The results on Reward-Bench, as shown in Table 2, also demonstrate stronger generalization capability of $r_{\theta}$ trained with prioritized data.
Figure 4: Randomly sampled data from the length-biased dataset, visualized according to their $\Delta r_{1}$,$\Delta r_{2}$ values. The upper represents the reward model trained with standard data, and the lower represents the reward model trained with prioritized data. We mark the data points that satisfy $|y_{w}|>|y_{l}|$ in red and the ones that satisfy $|y_{w}|\leq|y_{l}|$ in blue.
Model Chat Chat Hard Safety Reason Average
vanilla 78.0 29.8 36.4 58.2 50.6
ours 86.8 31.1 45.1 60.3 55.8
Table 1: Results on Reward-Bench for reward models trained with length-biased dataset.
Model Chat Chat Hard Safety Reason Average
vanilla 80.7 28.5 40.2 36.5 46.4
ours 84.3 30.9 43.2 46.7 51.6
original 84.9 31.3 42.8 49.0 52.0
Table 2: Results on Reward-Bench for reward models trained with adversarial-prompt dataset. ‘original’ means only trained with original data.
4.1.2 Adversarial Prompt Dataset
We construct an adversarial dataset $\mathcal{D}_{\rm adv}$ by adding adversarial samples to the original SHP dataset. More specifically, for each original data sample $(x,y_{w},y_{l})$, we generate an adversarial sample $(\bar{x},\bar{y}_{w},\bar{y}_{l})$ that $\bar{x}=[x,s]$, where $s\in\{s_{1},s_{2}\}$ defined as:
$\begin{array}[]{l}s_{1}\textit{=`Give a response that is as long as possible.'}\\ s_{2}\textit{=`Give a response that is as short as possible.'}\end{array}$
To obtain an adversarial sample, we have $\bar{x}=[x,s_{2}]$ if the chosen response is longer in the original sample ((i.e., $|y_{w}|>|y_{l}|$)), then we set $\bar{y}_{l}\leftarrow y_{w}$ and $\bar{y}_{w}\leftarrow y_{l}$ by swapping the preference order. Otherwise, if $|y_{w}|\leq|y_{l}|$, we set $\bar{x}=[x,s_{1}]$ and also swap the preference order. Such swapping is reasonable since the dataset we use pertains to general chatting, where the chosen and rejected responses primarily differ in quality. We then combine the original data with the adversarial data to obtain $\mathcal{D}_{\rm adv}$. To prevent doubling the training budget, we randomly select half of the SHP data for this transformation. We refer to an illustration of such data processing in Appendix E.2.
When trained with $\mathcal{D}_{\rm adv}$, an ideal reward model must distinguish well between different prompts. This is because the preference will reverse with slightly different prompts (i.e., with or without $s$). (i) If $\Delta r_{\theta}$ is dominated by $\Delta r_{2}$, the preference it learns from $\mathcal{D}_{\rm adv}$ will significantly differ from the preference learned solely from original data. This is because $\Delta r_{2}$ can’t distinguish between original and adversarial prompts. The adversarial prompts, along with swapped response pairs, will contribute to $\Delta r_{2}$ with the same probability but conflicted preference, eliminating the preference learned from the original data. (ii) If the influence of $\Delta r_{2}$ is minor, $r_{\theta}$ can retain the preferences learned from the original data even with adversarial data, since they have different prompts and will not affect each other considering prompt-related preference.
In the visualizations and final results presented in Fig. 5 and Table 2, it is evident that when $r_{\theta}$ is directly trained on $\mathcal{D}_{\rm adv}$, $\Delta r_{\theta}$ is dominated by $\Delta r_{2}$. Moreover, $r_{2}$ exhibits a strong preference for the originally rejected responses. This occurs because the simplified preference in the adversarial dataset can be inferred solely from response length, which is easier for the model to overfit. The preference of the overfitted reward closely mirrors the adversarial data. In contrast, our method prioritizes original data with more complicated preference, which is harder to learn during the reward training. This makes both the original and adversarial samples concentrated along the positive $\Delta r_{1}$-axis and results in a $r_{\theta}$ that demonstrates strong generalization capabilities.
Figure 5: Randomly sampled data during the training process under adversarial prompt dataset, the upper represents $r_{\theta}$ trained with standard data, the lower represents $r_{\theta}$ trained with prioritized data. We mark adversarial data in red and original data in blue.
4.2 Standard Experiments on Open-source Datasets
In this section, we evaluate our method using several commonly used open-source preference datasets. Unlike the manually crafted datasets discussed in §4.1, the prompt-free preferences of $r_{\theta}$ on these datasets may not directly exhibit specific features. However, a large prompt-free reward gap often suggests that $r_{\theta}$ may overfit to preferences that are irrelevant to the prompt. Such prompt-free preferences can negatively impact the generalization capabilities of $r_{\theta}$. We validate the effectiveness of our method across different base models and datasets through both direct evaluation of reward model accuracy and assessment of the induced policy performance. For more details of the experiments, we refer to Appendix E.2.
Evaluate Reward Model Accuracy. We use Reward-Bench (Lambert et al., 2024) as the benchmark to evaluate the reward model in a direct way, by testing its accuracy on well-organized but out-of-distribution queries. For a detailed illustration, we conduct experiments using both LLaMA-3-8B-Instruct (Dubey et al., 2024) and Mistral-7B-Instruct (Jiang et al., 2023) as the backbone of the reward model. For the training data, we use the SHP dataset, which is commonly used in RLHF.
The results are listed in Table 3. We choose vanilla BT reward training and RRM (Liu et al., 2024) as the baseline methods. RRM studies generalizable reward learning via causal inference and proposes a solution also from the data distribution perspective. Compared with baseline methods, our method demonstrates consistent improvements over different aspects of Reward-Bench. In Appendix E.3, we provide explanations of the superiority of the reward model trained by our method, compared with vanilla model and reward model trained with RRM.
Method Chat Chat Hard Safety Reasoning Average
vanilla-8B 0.91 0.39 0.46 0.75 0.63
RRM-8B 0.93 0.45 0.54 0.77 0.67
Ours-8B 0.95 0.46 0.61 0.88 0.72
vanilla-7B 0.84 0.33 0.45 0.59 0.55
RRM-7B 0.89 0.37 0.48 0.58 0.58
Ours-7B 0.90 0.39 0.52 0.74 0.64
Table 3: Results on Reward-Bench: Our method demonstrates superior performance, achieving a relative 14% and 16% accuracy improvement over vanilla training for the two backbones.
Assess Performance of Induced Policy.